# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v1

@@ Abstract
Diffusion models have demonstrated remarkable success in dense prediction problems, which aims to model per-pixel relationship between RGB images and dense signal maps, thanks to their ability to effectively capture complex data distributions. However, initiating the reverse sampling trajectory from uninformative noise prior introduces limitations such as degraded performance and slow inference speed. In this work, we propose DPBridge, a generative framework that formulates dense prediction tasks as image-conditioned generation problems and establishes a direct mapping between input image and its corresponding dense map based on fully-tractable diffusion bridge process. This approach addresses aforementioned limitations in conventional diffusion-based solutions. In addition, we introduce finetuning strategies to adapt our model from pretrained image diffusion backbone, leveraging its rich visual prior knowledge to facilitate both efficient training and robust generalization ability. Experimental results shows that our DPBridge can achieve competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, highlighting its effectiveness and adaptability.
Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing a stochastic transport between an RGB image and its corresponding dense map in the latent space based on a tractable diffusion bridge process. In addition, we introduce finetuning strategies to adapt large pretrained diffusion backbone to construct our bridge, leveraging its rich prior knowledge to enable both efficient training and robust generalization.
@@ 1 Introduction
Diffusion models recently emerge as a family of powerful generative models that are able to capture complex data distributions through a sequence of forward and reverse processes [26, 46]. They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image synthesis and video generation [17, 46, 27], controlled editing [61, 5, 50], human pose estimation [29, 30, 28], inverse problem solving [11, 53, 40].
Despite such success, diffusion models still face limitations that hinder broader applications, especially in discriminative tasks like dense prediction. One major limitation stems from the fact that standard diffusion reverse process always starts from Gaussian noise, which contains little information about the target distribution. This will lead to degraded performance or prolonged convergence times. Current diffusion-based approaches formulate this type of tasks as image-conditioned generation problems and adopt measures like incorporating external conditions to the model [47, 61, 25], modifying the generation pipeline [41, 54], or adding extra guidance terms during sampling [11, 53] to deal with this issue. In this way they hope to better model the conditional generation process.
Instead of starting from noise, an feasible alternative is to leverage the inherent problem structure. Since the goal of dense prediction tasks is to model the probability transport between image and dense signal map distributions, it’s a rather straightforward idea to directly start the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior compared to noise. This motivates the exploration of a new family of computational frameworks, i.e. diffusion bridge models [39, 63, 36, 37]. They are based on a special type of diffusion process conditioned on given endpoints, thus enabling optimal transport between two distributions and generalizing beyond Gaussian priors.
@@
Inspired by these intuition, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model adapted from pretrained image diffusion backbone to handle general dense prediction tasks though the image-conditioned generation pipeline. In essence, we construct diffusion bridge process between the latent distributions of images and dense signal maps, with all intermediate samples along the trajectory being convex combinations of the paired data. The generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from input images rather than pure noise, enabling faster inference speed as well as better generation quality.
@@
In addition, to better adapt pretrained diffusion backbone to our DPBridge construction, we also propose novel finetuning strategies to allow for more effective knowledge transfer. In this way our model is augmented with rich visual prior knowledge, which not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence. Fig.(1) is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e. depth estimation, surface normal prediction, and semantic segmentation, proving that our framework is theoretically universal and can be instantiated to different scenarios.
@@
  • We propose DPBridge framework to deal with dense prediction tasks with generative pipeline by formulating them as instances of latent diffusion bridges.
@@ 2.1 Dense Prediction
Dense prediction refers to the type of problems whose outputs match the spatial dimensions of input data, often with the goal of assigning a value to each individual pixel. Typical examples include depth estimation, surface normal prediction, semantic segmentation, optical flow, etc. Common solutions can be categorized in discriminative and generative manner. Discriminative approaches focus on architecture design and introduce techniques like skip connections, dilated convolutions, multi-scale feature aggregation, and attention mechanisms [9, 44, 34] to better capture both local and global context for improved prediction accuracy. Generative methods, on the other hand, offer an alternative paradigm to tackle these problems through distribution modeling. Some works reuse or merge prior features from pretrained vision foundation models [4, 23, 62]. Others formulate these tasks as visual-condition guided generation process which refines the desired dense map iteratively [18, 32, 24].
2.2 Diffusion Bridge Models
Stochastic bridge models are a common tool in probability theory which describe the evolution of stochastic processes with given starting/ending points. When combined with diffusion models, they provide a new perspective to translate between different distributions and model conditional probabilities without the need for prior knowledge, therefore have been gaining increasing attention in the context of generative modeling in recent year [39, 38, 16]. They demonstrate the advantages of data-to-data generation paradigm over noise-to-data one on various tasks, including image-to-image translation [63, 37, 36], image restoration [59, 13], text-to-speech synthesis [10], etc. However, to the best of our knowledge, no prior work has explored leveraging diffusion bridge models on dense prediction tasks to facilitate the distribution transport between images and dense signal maps, despite their intuitive structural closeness. Our work aims to fill in this research gap.
@@ 3.1 Diffusion Process
where $\mathbf{w}_{t}$ is a standard Wiener process, $f:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ and $g:[0,T]\rightarrow\mathbb{R}$ are the drift and diffusion coefficients that describe the evolution of this stochastic process. The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\sigma_{t}$ depends on the design choice of drift and diffusion coefficients. It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$
@@ 3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that guarantee to arrive at a prescribed final state $T$. One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints. We can use Doob’s h-transform technique [42, 48] to modify a reference process $\mathbb{Q}$ defined in Eqn.(1) to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens. Therefore the conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ follows the law of
@@
where $h(\mathbf{z}_{t},t,\mathbf{x},T)=\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{T}=\mathbf{x}|\mathbf{z}_{t})$. Intuitively, the additional drift term $h(\mathbf{z}_{t},t,\mathbf{x},T)$ plays the role of steering $\mathbf{z}_{t}$ towards the specified target point $\mathbf{z}_{T}=\mathbf{x}$, and it is ensured that $\mathbb{Q}^{\mathbf{x}}$ invariably passes $\mathbf{x}$ at timestep $T$ with probability 1 for any initial state $\mathbf{z}_{0}$.
@@ 4.1 Problem Formulation
We formulate different dense prediction tasks like depth estimation, surface normal prediction, semantic segmentation etc., as image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g. segmentation, depth, surface normal), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image. In the following sections, we’ll first introduce the motivation of utilizing diffusion bridge models to handle dense prediction problems. Then, we’ll elaborate the details of our proposed DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
Figure 2: Overview of our training pipeline, which is conducted in the latent space based on diffusion bridge theory. To overcome the discrepancy with standard diffusion process and facilitate finetuning, we derive the maximum likelihood training objective of diffusion bridge process, propose a distribution aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
4.2 Motivation: Bridge VS Diffusion
As introduced previously, the incompatibility that employ diffusion models to dense prediction task, which is essentially a image-conditioned generation problem, lies in its underlying iterative denoising process starting from Gaussian noise. In such cases, the distance between prior Gaussian distribution and data distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms to converge to target distributions, at the risk of degraded performance and slow inference speed.
Bridge models, on the other hand, directly formulate generative processes as conditional probability paths. In this scenario, the boundary distribution is not restricted to Gaussian noise, but instead can be chosen as more informative priors, such as the condition data distribution. By establishing direct connections between different distributions, this formulation provides a more natural and efficient way to model probability transport compared with previous efforts. For example, it reduce the dependence on manually designed post-hoc guidance [11, 53, 12] or customized generation pipeline [41]. In addition, through implicitly integrates condition as part of the learned process, it provides a more flexible way to model intricacies between data and conditions without relying on explicit mechanisms like concatenation, attention [47, 61, 25], whose inductive bias may harm the overall performance when dealing with complex circumstances. Considering these advantages, we advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
@@ 4.3 Dense Prediction Bridge
Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Sec.3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more severe, it lacks the ability to harness rich scene understanding knowledge inherent in the vision foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks. Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between original diffusion bridge process and standard diffusion process, aiming for more efficient knowledge transfer. Fig.2 gives an overview of the proposed finetuning pipeline, and we’ll explain the design of each component next.
@@ 4.4 Latent Space Implementation
Instead of directly learning in image pixel space, we follow common practices to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the target dense maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the boundary states. In this way the diffusion bridge process is henceforth learned in the latent space. To retrieve original signal map we can use the decoder to decode the predicted latents $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. In later part of this paper, with little abuse of notation, all derivations are conducted in the latent space.
@@ 4.5.1 Maximum Likelihood Loss
In previous diffusion bridge models, widely-used techniques for learning the reverse process that corresponds to Eqn.(2) is to follow the theory of [1], use a neural network to approximate the score function and apply score matching technique. We notice that analytical form of diffusion bridge’s reverse transition kernel is directly obtainable, therefore the usage of maximum likelihood training can also be applied, which is performed by optimizing the Evidence Lower Bound (ELBO). It yields a more stable loss function compared with score matching and better aligns with the pretrained backbone. Given the fixed terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
$\text{ELBO}=&\mathbb{E}_{p(\mathbf{z}_{0})}\Big{[}\mathbb{E}_{p(\mathbf{z}_{1}\mid\mathbf{z}_{0})}\big{[}\log p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{1},\mathbf{z}_{T})\big{]}-\\ \sum_{t=2}^{T}\mathbb{E}_{p(\mathbf{z}_{t}|\mathbf{z}_{0})}&\big{[}\mathbb{D}_{KL}\big{(}q(\mathbf{z}_{t-1}\mid\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\big{)}\big{]}\Big{]}\end{aligned}$
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process and $\mathbb{D}_{KL}$ is the KL divergence. Utilize Bayes’ theorem and Markov chain property, we obtain:
@@
Similar to the strategy in DDPM [26], we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as:

### v2

@@ Abstract
Diffusion models have shown remarkable capabilities in modeling complex data distributions by transforming noise into structured data through stochastic processes. However, when applied to dense prediction tasks whose goal is to capture per-pixel relationships between RGB images and dense signal maps, starting the sampling process from an uninformative Gaussian noise often leads to inefficient sampling and long latency. To overcome these challenges, we propose DPBridge, a generative framework that establishes direct mapping between input RGB images and dense signal maps based on a tractable bridge process. Furthermore, we introduce finetuning strategies to leverage a pretrained large-scale image diffusion backbone, enjoying its rich visual prior knowledge to enable both efficient training and robust generalization. Experiments show that DPBridge achieves competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, validating its effectiveness and adaptability.
Machine Learning, ICML
Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing conditional probability transport between RGB images and corresponding dense maps in the latent space based on a tractable diffusion bridge process. In addition, we introduce finetuning strategies to adapt large pretrained diffusion backbone to construct our bridge, leveraging its rich prior knowledge to enable both efficient training and robust generalization.
@@ 1 Introduction
Dense prediction refers to the type of problems whose primary goal is to assign a value at every spatial location in the input data, generating output signal maps that possess same structure and size as inputs. It incorporates various typical tasks, including depth estimation (Yang et al., 2024; Ranftl et al., 2021; Eftekhar et al., 2021), surface normal prediction (Bae & Davison, 2024; Hu et al., 2024; Bae et al., 2021), semantic segmentation (Chen, 2017; Cheng et al., 2022; Wang et al., 2023; Weber et al., 2021), optical flow (Teed & Deng, 2020; Huang et al., 2022; Weinzaepfel et al., 2023), etc.
Recently, image diffusion models have been applied to dense prediction tasks by formulating them as conditional generation problems. Their ability to capture complex data distributions makes them effective priors, leading to highly favorable results (Ke et al., 2024; Garcia et al., 2024; Fu et al., 2025). Following a common practice, these methods diffuse signal maps into pure noise and subsequently leverage input RGB images as conditions to guide the generation process. Nonetheless, there still exists limitations in diffusion-based solutions, mostly stem from the underlying modeling mechanism. The denoising process of a standard diffusion model always starts from Gaussian noise, which contains little information about the target signal distribution. In such cases, the distance between prior Gaussian distribution and target distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms, at the risk of degraded performance and prolonged convergence time.
Instead of starting from noise, one feasible alternative is to directly initiate the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior. This leads to the exploration of a new family of computational frameworks, i.e. diffusion bridge models (Liu et al., 2022; Zhou et al., 2023; Li et al., 2023; Liu et al., 2023). They are based on stochastic process with fixed starting and ending points and are adept at modeling conditional probability paths. By establishing direct connections between different distributions, this line of framework provides a more natural and efficient way to model probability transport compared with previous efforts by reducing the dependence on manually designed post-hoc guidance (Chung et al., 2022a; Song et al., 2023; Chung et al., 2022b) or customized generation pipeline (Meng et al., 2021). We therefore advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
@@
Motivated by this, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model to handle general dense prediction tasks through generative pipeline. In essence, we modify the standard diffusion process to directly bridge the latents of RGB images and corresponding dense signal maps, and enforce the model to learn this process through maximum likelihood estimation. In this way the generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from RGB images rather than pure noise, enabling faster inference speed as well as better generation quality.
@@
Furthermore, we introduce novel finetuning strategies to better adapt pretrained diffusion backbone to DPBridge construction. The equipped rich visual prior not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence. Figure 1 is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e. depth estimation, surface normal prediction, and semantic segmentation, proving that our framework is theoretically universal and can be instantiated to different scenarios.
@@
  • We propose the DPBridge framework to address dense prediction tasks within a generative pipeline by formulating them as instances of latent diffusion bridges.
@@ 2.1 Dense Prediction
Dense predictions can be characterized as a conditional distribution modeling task and common solutions can be categorized into discriminative and generative ones. Discriminative methods directly map input data to structured outputs in a feed-forward manner. (Ranftl et al., 2021) introduces attention mechanism and transformer architecture to better capture both local and global context feature. (Yang et al., 2024) explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness. (Bae & Davison, 2024) incorporates per-pixel ray directions and models the relative rotations between neighboring pixels. Recently, diffusion-based generative models have been applied in this field and showed remarkable performance. (Ke et al., 2024) finetunes pretrained diffusion model, leveraging the rich visual knowledge within to enhance depth estimation. (Fu et al., 2025) takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions. (Gui et al., 2024) exploits flow matching to facilitate the distribution shift between images and depth maps. Our work follow conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in iterative denoising manner.
2.2 Diffusion and Bridge Models
Diffusion models (Ho et al., 2020; Song et al., 2020b) are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes. They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis (Rombach et al., 2022; Ho et al., 2022), controlled editing (Zhang et al., 2023; Batzolis et al., 2021), inverse problem solving (Chung et al., 2022a; Song et al., 2023), 3D assets generation (Xiang et al., 2024; Zhang et al., 2024), and human-centered applications (Ji et al., 2024b, a; Ji & Li, 2024).
Diffusion bridge models have gained increasing attention in generative modeling in recent years (Liu et al., 2022; Zhou et al., 2023; Liu & Wu, 2023; De Bortoli et al., 2021). They offer a novel perspective for translating between different distributions and enable a more effective data-to-data generation paradigm. Recent works have explored different types of bridge processes and their applications. (Liu et al., 2023; Li et al., 2023; Yue et al., 2023) develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution. (Chung et al., 2024) proposes a modified inference procedure that enforces data consistency. (Chen et al., 2023) introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency. To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness. Our work aims to address this research gap.
@@ 3.1 Diffusion Process
where $\mathbf{w}_{t}$ is a standard Wiener process, $f:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ and $g:[0,T]\rightarrow\mathbb{R}$ are the drift and diffusion coefficients that describe the evolution of this stochastic process. The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\sigma_{t}$ depends on the design choice of drift and diffusion coefficients. It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$.
@@ 3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that is conditioned to end at a specified terminal point at timestep $T$. One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints. We can use Doob’s h-transform technique (Oksendal, 2013; Särkkä & Solin, 2019) to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value. The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
@@
where $h(\mathbf{z}_{t},t,\mathbf{x},T)=\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{T}=\mathbf{x}|\mathbf{z}_{t})$. Intuitively, the additional drift term $h(\mathbf{z}_{t},t,\mathbf{x},T)$ plays the role of steering $\mathbf{z}_{t}$ towards the target point $\mathbf{z}_{T}=\mathbf{x}$, and it is ensured that $\mathbb{Q}^{\mathbf{x}}$ invariably passes $\mathbf{x}$ at timestep $T$ with probability 1 for any initial point $\mathbf{z}_{0}$.
@@ 4.1 Problem Formulation
We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g. depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image. In the following sections, we’ll elaborate the details of DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
Figure 2: Overview of our training pipeline. To overcome the discrepancy with standard diffusion process and facilitate finetuning, we derive the maximum likelihood training objective of diffusion bridge process, propose a distribution aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
@@ 4.2 Dense Prediction Bridge
Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Section 3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more, it lacks the ability to harness rich visual prior knowledge inherent in the foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks. Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between bridge process and standard diffusion process, aiming for more efficient knowledge transfer. Figure 2 gives an overview of the proposed finetuning pipeline, and next we’ll explain the design of each component.
@@ 4.3 Latent Space Implementation
Instead of directly learning in image pixel space, we follow common practice to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the dense signal maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the initial and terminal points. In this way the diffusion bridge process is henceforth learned in the latent space. To retrieve original signal map we can use the decoder to decode the predicted latents $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. In later part of this paper, with little abuse of notation, all derivations are conducted in the latent space.
@@ 4.4.1 Maximum Likelihood Loss
Widely-used techniques for learning the reverse process that corresponds to Equation 2 is to follow the theory of (Anderson, 1982), use a neural network to approximate the score function and apply score matching. We notice that analytical form of diffusion bridge’s reverse transition kernel is directly obtainable, therefore the usage of maximum likelihood training can also be applied, which is performed by optimizing the Evidence Lower Bound (ELBO). It yields a more stable loss function compared with score matching and better aligns with the pretrained backbone. Given the specified terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
[…]
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process and $\mathbb{D}_{KL}$ is the KL divergence. Utilizing Bayes’ theorem and Markov chain property, we can obtain:
@@
Similar to the strategy in DDPM (Ho et al., 2020), we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as: